# Programmable k-local Ising Machines and all‑optical Kolmogorov-Arnold Networks on Photonic Platforms

### v1

@@ Abstract
We unify k-local Ising optimization and optical KAN function learning on a single photonic platform, establishing a critical convergence point in optical computing that enables interleaved discrete-continuous workflows. We introduce a single spacial light modulator (SLM)-centric primitive that realizes, in one stroke, all-optical k-local Ising interactions and fully optical Kolmogorov-Arnold network (KAN) layers. The central idea is to convert structural nonlinearity of a nominally linear photonic scatterer into a per-window computational resource by adding one relay pass through the same spatial light modulator. A folded 4f relay reimages the first Fourier plane onto the SLM so that each chosen spin clique or ridge channel occupies a disjoint window with its own second-pass phase patch. Propagation remains linear in the optical field, yet the measured intensity in each window becomes a freely programmable polynomial of the clique sum or projection amplitude. This yields native, per-clique k-local couplings without nonlinear media and, in parallel, the many independent univariate nonlinearities required by KAN layers, all with in-situ physical gradients for training using two-frame (forward and adjoint) physical gradients. We outline implementation on spatial photonic Ising machines, injection-locked VCSEL arrays, and the Microsoft analog optical computers. In all cases the hardware change is one extra lens and a fold (or an on-chip 4f loop), enabling a minimal overhead, massively parallel route to high-order optical Ising optimization and trainable, all-optical KAN processing.
@@ I Introduction
Ising machines are physical simulators that seek low-energy configurations of an Ising-type cost function by exploiting analog optical, electronic, or quantum dynamics. They have emerged as high-throughput engines for quadratic unconstrained binary optimization (QUBO) and for sampling from near-Boltzmann distributions [1]. Realizations now include free‑space SLM‑based photonic Ising machines, polaritonic and photonic Ising machines, large‑scale coherent Ising machines and integrated DOPO networks, as well as electronic annealers and Ising‑inspired engines (simulated bifurcation, analog iterative machines) and spintronic p‑bit computers; see, e.g., [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15]. Optical realizations are particularly attractive because diffraction and interference deliver native, reconfigurable all-to-all linear coupling with negligible latency and excellent parallelism; in free-space spatial-photonic systems millions of spins can be encoded on a single spatial light modulator (SLM) aperture [8], while injection-locked vertical-cavity surface-emitting laser (VCSEL) arrays achieve sub-nanosecond state updates at rack-scale densities [16]. These attributes make optics a natural substrate for both native kk‑local interactions and for the per‑ridge nonlinearities required by Kolmogorov-Arnold networks (KANs), motivating the unified approach developed below.
@@
KAN-type architectures have universal-approximation guarantees (e.g., P1‑KAN; KKANs). Formal comparisons proving a global, orders‑of‑magnitude parameter reduction over multilayer perceptrons (MPL) for universal approximation do not exist to our knowledge. Empirically, there are tasks (e.g., Poisson PDE) where a small KAN matches or exceeds the accuracy of a much larger MLP with roughly 100× fewer parameters [29]. The first photonic KAN implementation demonstrated $65\times$ reduction in energy consumption and $50\times$ improvement in latency compared to traditional photonic neural networks, validating the potential for optical KAN realizations [30]. KANs have been deployed commercially for credit‑risk modeling and fraud detection [31]. Equation (3) is natural for optical implementation: the projections are realized by beam-splitting or diffraction, and each $\Phi_{m}$ is an independent, univariate nonlinearity. Critically, an all-optical KAN therefore requires many independent, per-ridge, programmable univariate nonlinearities. With this context, we now summarize the operational status and problem reach of the three optical lines of attack most relevant to this work.
@@
Microsoft’s analog optical computer (AOC) performs the matrix–vector multiplication in optics and realizes all remaining primitives of the fixed‑point iteration in analog electronics [12]. The state $s_{t}\in\mathbb{R}^{N}$ at time $t$ is encoded as intensities of a micro light‑emitting diode (microLED) array and multiplied optically by an SLM that stores the weight matrix $W\in\mathbb{R}^{N\times N}$. Three‑dimensional (3D) fan‑in/fan‑out free‑space optics distribute light from each microLED across an SLM row for element‑wise modulation, and the signals are summed column‑wise on a photodetector (PD) array to produce the optical matrix-vector product (by optical vector-matrix multiplication, OVMM). Signed weights are implemented by loading the positive and negative parts of $W$ onto two SLMs and combining the corresponding PD outputs by differential analog readout. The resulting channel voltages undergo an element‑wise analog nonlinearity $f:\mathbb{R}^{N}\to\mathbb{R}^{N}$ and configurable per‑channel gains that implement an annealing schedule $\alpha(t)\in\mathbb{R}$ and a momentum coefficient $\gamma\in\mathbb{R}$. One discrete iteration (corresponding to a single optical–electronic round trip) is

### v2

@@ Abstract
Photonic computing promises energy-efficient acceleration for optimization and learning, yet discrete combinatorial search and continuous function approximation have largely required distinct devices and control stacks. Here we unify k-local Ising optimization and optical Kolmogorov-Arnold network (KAN) learning on a single photonic platform, establishing a critical convergence point in optical computing. We introduce an SLM-centric primitive that realizes, in one stroke, all-optical k-local Ising interactions and fully optical KAN layers. The key idea is to convert the structural nonlinearity of a nominally linear scatterer into a per-window computational resource by adding a single relay pass through the same spatial light modulator: a folded 4f relay re-images the first Fourier plane onto the SLM so that each selected clique or channel occupies a disjoint window with its own second pass phase patch. Propagation remains linear in the optical field, yet the measured intensity in each window becomes a freely programmable polynomial of the clique sum or projection amplitude. This yields native, per clique k-local couplings without nonlinear media and, in parallel, the many independent univariate nonlinearities required by KAN layers, all trainable with in-situ physical gradients using two frames (forward and adjoint). We outline implementations on spatial photonic Ising machines, injection-locked vertical cavity surface emitting laser (VCSEL) arrays, and Microsoft analog optical computers; in all cases the hardware change is one extra lens and a fold (or an on-chip 4f loop), enabling a minimal overhead, massively parallel route to high-order Ising optimization and trainable, all-optical KAN processing on one platform.
@@ I Introduction
Ising machines are physical simulators that seek low-energy configurations of an Ising-type cost function by exploiting analog optical, electronic, or quantum dynamics. They have emerged as high-throughput engines for quadratic unconstrained binary optimization (QUBO) and for sampling from near-Boltzmann distributions [1]. Realizations now include free‑space SLM‑based photonic Ising machines, polaritonic and photonic Ising machines, large‑scale coherent Ising machines and integrated DOPO networks, as well as electronic annealers and Ising‑inspired engines (simulated bifurcation, analog iterative machines) and spintronic p‑bit computers; see, e.g., [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16]. Optical realizations are particularly attractive because diffraction and interference deliver native, reconfigurable all-to-all linear coupling with negligible latency and excellent parallelism; in free-space spatial-photonic systems millions of spins can be encoded on a single spatial light modulator (SLM) aperture [8], while injection-locked vertical-cavity surface-emitting laser (VCSEL) arrays achieve sub-nanosecond state updates at rack-scale densities [17]. These attributes make optics a natural substrate for both native kk‑local interactions and for the per‑ridge nonlinearities required by Kolmogorov-Arnold networks (KANs), motivating the unified approach developed below.
@@
KAN-type architectures have universal-approximation guarantees. Formal comparisons proving a global, orders‑of‑magnitude parameter reduction over multilayer perceptrons (MPL) for universal approximation do not exist to our knowledge. Empirically, there are tasks (e.g., Poisson PDE) where a small KAN matches or exceeds the accuracy of a much larger MLP with roughly 100× fewer parameters [30]. The first photonic KAN implementation demonstrated $65\times$ reduction in energy consumption and $50\times$ improvement in latency compared to traditional photonic neural networks, validating the potential for optical KAN realizations [31]. KANs have been deployed commercially for credit‑risk modeling and fraud detection [32]. Equation (3) is natural for optical implementation: the projections are realized by beam-splitting or diffraction, and each $\Phi_{m}$ is an independent, univariate nonlinearity. Critically, an all-optical KAN therefore requires many independent, per-ridge, programmable univariate nonlinearities. With this context, we now summarize the operational status and problem reach of the three optical lines of attack most relevant to this work.
@@
Microsoft’s analog optical computer (AOC) performs the matrix–vector multiplication in optics and realizes all remaining primitives of the fixed‑point iteration in analog electronics [12, 13]. The state $s_{t}\in\mathbb{R}^{N}$ at time $t$ is encoded as intensities of a micro light‑emitting diode (microLED) array and multiplied optically by an SLM that stores the weight matrix $W\in\mathbb{R}^{N\times N}$. Three‑dimensional (3D) fan‑in/fan‑out free‑space optics distribute light from each microLED across an SLM row for element‑wise modulation, and the signals are summed column‑wise on a photodetector (PD) array to produce the optical matrix-vector product (by optical vector-matrix multiplication, OVMM). Signed weights are implemented by loading the positive and negative parts of $W$ onto two SLMs and combining the corresponding PD outputs by differential analog readout. The resulting channel voltages undergo an element‑wise analog nonlinearity $f:\mathbb{R}^{N}\to\mathbb{R}^{N}$ and configurable per‑channel gains that implement an annealing schedule $\alpha(t)\in\mathbb{R}$ and a momentum coefficient $\gamma\in\mathbb{R}$. One discrete iteration (corresponding to a single optical–electronic round trip) is